OpenStudy (anonymous):

2y-4/4=3,find the value of y

OpenStudy (lgbasallote):

cross multiply... 2y - 4 = 12 2y = 12 + 4 2y = 16 y = 8

OpenStudy (mayliz):

y=2

OpenStudy (anonymous):

\[2y-4 = 4\times3\] \[2y = 16\] \[y=8\]

OpenStudy (mayliz):

2y-1=3 2y=4 y=2

OpenStudy (anonymous):

If there is no parentheses 2y - 4/4 = 3 2y - 1 = 3 2y = 4 y = 4/2 = 2 If there is parentheses (2y - 4)/4 = 3 2y - 4 = 12 2y = 16 y = 16/2 = 8